The parallel sides of a trapezium are in the ratio 2:3 and the area of the trapezium is 125 cm square. The distance between the parallel lines is 10cm. Find the length of the parallel sides of the trapezium.

Answer by fractalier(6550)

The formula for the area of a trapezoid is
A = (h/2)(b1 + b2)
Here h = 10, b1 = 2/3(b2), and A = 125...now plug in
125 = (10/2)(b + (2/3)b)
25 = (5/3)b
b = 15 cm
2/3 b = 10 cm